A lottery has 52 numbers. In how many different ways can six of the numbers be selected? (Assume that order of selection is not important.)
what principle is used in this question?

Answer by user_dude2008(1862)

Use the combination formula

52 C 6 = 52!/(6!(52-6)!)=52!/(6!46!)=(52*51*50*49*48*47)/(6*5*4*3*2*1)=20358520

Answer: There are 20,358,520 different ways to select 6 numbers (from 52 possible)